OpenStudy (anonymous):

find the equation of a line passing through (-4,3) and (1,2)

OpenStudy (anonymous):

First find slope

OpenStudy (anonymous):

m=(2-3)/(1+4)=-1/5

OpenStudy (aroub):

ok first thing u find the slope 2-3/1+4 -1/5 then u choose any of the points u have ill choose (1,2) to find the y-intercept y=mx+b 2=-1/5(1)+b , 2=-1/5+b , 10/5+1/5 =b , 11/5 =b the equation of the line is y=-1/5 +11/5

OpenStudy (aroub):

y=-1/5x +11/5 **

OpenStudy (anonymous):

Now use \[y-y _{1}=m(x-x _{1})\]